OpenStudy (anonymous):

how do I finish this If π < t < 3π/2 then π/2 < t/2 < 3π/4; for such values of t, cos(t/2) < 0 and sin(t/2) > 0 From this and other trig identities, cos(t/2) = -√[(1 + cos(t))/2] sin(t/2) = √[(1 - cos(t))/2] how do I finish this If π < t < 3π/2 then π/2 < t/2 < 3π/4; for such values of t, cos(t/2) < 0 and sin(t/2) > 0 From this and other trig identities, cos(t/2) = -√[(1 + cos(t))/2] sin(t/2) = √[(1 - cos(t))/2] @Mathematics

OpenStudy (anonymous):

So what are you suppose to do on this?? Im kinda lost

OpenStudy (anonymous):

this is the original question

OpenStudy (anonymous):

If cos(t)=−3/4 where π<t<3π/2, find the values of the following trigonometric functions? cos(2t) = sin(2t)= cos(t/2)= sin(t/2)=

OpenStudy (anonymous):

i have found the first two to be cos(2t) = 2 cos²(t) - 1 = 2(-¾)² - 1 = 1/8 sin(t) = -√(1 - cos²(t)) = -√(1 - (-¾)²) = -√7 / 4 sin(2t) = 2 sin(t) cos(t) = 2(-√7/4)(-¾) = 3 √7 / 8

OpenStudy (anonymous):

and i started the remaining two but dont know how to finish it

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\cos(t/2) = \sqrt{\frac{1+\cos(x)}{2}}\]

OpenStudy (anonymous):

that cos(x) should be cos(t) in this case.

OpenStudy (anonymous):

\[\cos(t/2) = \sqrt{\frac{1+(-\frac{3}{4})}{2}}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

im having extreme difficulties entering this into my homework software lol

OpenStudy (anonymous):

you should get \[\frac{\sqrt{8}}{8}\]

OpenStudy (anonymous):

it rejected that..uhhh